OpenStudy (bmoneythewizard):

Find limit as x approaches two of quantity x squared minus four divided by quantity x minus two. You must show your work or explain your work in words.

Nnesha (nnesha):

\[\huge\rm \lim_{x \rightarrow 2} \frac{ x^2-4 }{ x-2 }\] like this ???

Nnesha (nnesha):

direct substitution: if you substitute x=2 you will get 0 at the denominator therefore.. first simplify the fraction difference of squares \[\large\rm a^2-b^2 =(a+b)(a-b)\]

OpenStudy (evoker):

Of course for later problems if this wasn't available you could L'Hopital's rule

OpenStudy (karim728):